OpenStudy (anonymous):

What is the value of x in the equation 7x + 2y = 48 , when y = 3?

OpenStudy (igreen):

Plug in 3 for 'y': 7x + 2y = 48 7x + 2(3) = 48 2 * 3 = ?

OpenStudy (igreen):

Can you multiply 2 * 3?

OpenStudy (anonymous):

6

OpenStudy (igreen):

Yes, so we have: 7x + 6 = 48 Subtract 6 to both sides, can you do that?

OpenStudy (anonymous):

7x = 42

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

yes

OpenStudy (igreen):

Yep! Now divide 7 to both sides.

OpenStudy (anonymous):

so the answer is 6

OpenStudy (igreen):

Yep!

OpenStudy (igreen):

\(\color{lime}{\underbrace{\overbrace{\Huge\bigstar\star\ast\bigstar\star\ast\bigstar\star\ast\bigstar\star\ast\\\bigstar\star\ast\bigstar\star\ast\bigstar\star\ast\bigstar\star\ast\bigstar\star\ast\bigstar\star\ast\bigstar\star\ast\bigstar\star\ast\bigstar\star\ast\bigstar\star\ast\\\Large{~\succ}\bigstar\boxed{\sf Welcome~to~\color{#3dabea}{Open}\color{#9cc575}{Study}!!!}\bigstar\hspace-5px\prec\\\bigstar\star\ast\bigstar\star\ast\bigstar\star\ast\bigstar\star\ast\bigstar\star\ast\bigstar\star\ast\bigstar\star\ast\bigstar\star\ast\bigstar\star\ast\bigstar\star\ast\\\Huge{\bigstar\star\ast\bigstar\star\ast\bigstar\star\ast\bigstar\star\ast}}}}\) \(\small\sf\color{red}{Made~by~@nobrainiachere,~edited~by~@iGreen}\) You can give medals by clicking 'Best Response'.